# BayTTA: Uncertainty-aware medical image classification with optimized test-time augmentation using Bayesian model averaging

### v1

@@ 4 Experimental evaluations and discussion
4.1 Data used in evaluation
We conducted our experiments on three publicly available medical image datasets: skin cancer, breast cancer, and chest X-ray data, which are freely available on the Kaggle and Mendeley data repositories (see also Fig. 1). Subsequently, we evaluated the efficacy of the proposed BayTTA method using two well-known gene editing datasets, CRISPOR and GUIDE-seq, both generated through the CRISPR-Cas-9 technology ([50]). We adopted an effective one-hot encoding strategy, originally proposed by [10], to represent sgRNA and DNA sequence pairs as binary images from which off-target effects can be predicted by advanced DL models (see Fig. 4).
4.1.1 Medical image datasets
@@
Chest X-ray dataset is a large collection of X-ray images extracted from a publicly available medical image database [^3] https://easy.dans.knaw.nl/ui/datasets/id/easy-dataset:102451 ([33]). The dataset contains 5216 training images, among which 1341 are categorized as Normal and 3875 as Pneumonia samples. Additionally, the test dataset includes 624 images, among which 234 are classified as Normal and 390 as Pneumonia samples.
4.1.2 Gene editing datasets
CRISPOR database, organized and maintained by Maximilian Haeussler ([23]) aggregates widely-used public datasets aimed at quantifying on-target guide efficiency and detecting off-target cleavage sites [^4] http://crispor.tefor.net. The dataset we selected from this database (see also [10]) comprises $18,211$ black and white training images (each with $8\times 23$ pixel dimension), among which 18,112 are categorized as on-targets and 99 as off-targets. Additionally, the dataset includes 7806 testing images of the same dimension, 7763 of which are classified as on-targets and 43 as off-targets.
GUIDE-seq is one of the pioneering off-target data repositories, derived from the outcomes of the GUIDE-seq technique developed by [55]. It serves as an accurate framework for genome-wide identification of off-target effects. The sgRNAs used in GUIDE-seq target the following sites: VEGFA site 1, VEGFA site 2, VEGFA site 3, FANCF, HEK293 site 2, HEK293 site 3, and HEK293 site 4, wherein 28 off-targets with a minimum modification frequency of $0.1$ were identified (among 403 potential off-targets). This dataset consists of black and white images with $8\times 23$ pixel dimension. It comprises 309 training images, including 291 on-target and 18 off-target samples. Additionally, it comprises 133 testing images, including 121 on-target and 12 off-target samples.
Fig. 4: Examples of visualizing CRISPR-Cas9 sgRNA-DNA sequence pairs encoded onto 8×23 matrices, then transformed into black and white images from the (a) CRISPOR and (b) GUIDE-seq gene editing datasets, respectively ([10]). These images can be processed by neural networks to predict off-targets generated by CRISPR-Cas9 technology.
Table 1: Hyperparameter configuration of the pre-trained deep learning models (VGG-16, MobileNetV2, DenseNet201, ResNet152V2, and InceptionResNetV2) for the Skin Cancer, Breast Cancer, and Chest X-ray medical image datasets considered in our study.
@@ 4.1.2 Gene editing datasets
4.2 Implementation details and model settings
@@
Pre-trained CNN models. During the training phase of our first experiment, we employed five well-established convolutional networks (CNNs), namely VGG-16, MobileNetV2, DenseNet201, ResNet152V2, and InceptionResNetV2, as our feature extraction backbone, initializing weights through pre-training on the ImageNet dataset. The baseline versions of these models were initially used (without applying TTA or BayTTA), followed by their application in combination with TTA and BayTTA. All medical images were resized to the $224\times 224$ pixel size to ensure uniformity in model inputs. The hyperparameters, including the optimizer (Opt), the learning rate (LR), the batch size (BS), and the number of epochs for each of the five pre-trained models and each medical image dataset are summarized in Table 1. Additionally, we employed the early stopping technique to handle overfitting ([44]). During our network training, the data augmentation techniques were implemented using the following parameters: rotation range = 20, width shift range = 0.2, height shift range = 0.2, shear range = 0.2, zoom range = 0.2, horizontal flip, and vertical flip.
@@
Four pre-trained CNN models, VGG-16, MobileNetV2, DenseNet201, and ResNet152V2, as well as a custom CNN with five layers, comprising two convolutional and three fully connected dense layers, were used to predict off-targets in highly imbalanced gene editing datasets. The models were trained using the RMSprop optimizer, with a learning rate of $0.001$ and a batch size of $64$. To address potential overfitting issue a set of callback functions was used during training, specifically the ReduceLROnPlateau and EarlyStopping functions.
@@
State-of-the-art DL models intended for medical image analysis. To evaluate the performance of BayTTA, we also compared its performance against TTA and baseline, considering some state-of-the-art DL models recently used in medical image analysis, namely Attention Residual Learning (ARL) ([61]), COVID-19 ([43]), LoTeNet ([45]), IRv2+SA ([14]), and PCXRNet ([18]).
@@
IRv2+SA ([14]) explores the efficacy of soft attention in deep neural architectures. The core objective of this technique is to emphasize the significance of essential features, while mitigating the influence of noise-inducing ones. The ARL model, proposed by [61], is intended for classifying skin lesions in dermoscopy images. The model architecture includes several ARL blocks, a global average pooling layer, and a classification layer. Each ARL block employs a combination of residual learning and an innovative attention-learning mechanism to boost the capacity of discriminative representations. [43] proposed a novel model based on the DarkNet method designed for automated detection of COVID-19 cases using chest X-ray images to deliver precise diagnostic results in the framework of binary and multi-class classifications. PCXRNet ([18]) is an attention-driven convolutional neural network designed for pneumonia diagnosis based on chest X-ray image analysis. PCXRNet incorporates a novel Condensed Attention Module (CDSE) to harness the information within the feature map channels. The LoTeNet (Locally order-less Tensor Network) model ([45]) is based on the use of tensor networks, a crucial tool for physicists to analyze complex quantum many-body systems.
4.3 Experimental results
@@
Table 2: Comparison of the baseline CNN model accuracy (%) $\pm$ STD performance against the TTA and BayTTA versions on the skin cancer dataset. The highest accuracy per column is in bold. The asterisk (*) denotes the highest overall accuracy obtained by the models.
@@
Table 3: Comparison of the baseline CNN model accuracy (%) $\pm$ STD performance against the TTA and BayTTA versions on the breast cancer dataset. The highest accuracy per column is in bold. The asterisk (*) denotes the highest overall accuracy obtained by the models.
@@
Table 4: Comparison of the baseline CNN model accuracy (%) $\pm$ STD performance against the TTA and BayTTA versions on the chest X-ray dataset. The highest accuracy per column is in bold. The asterisk (*) denotes the highest overall accuracy obtained by the models.
@@
Table 5: Comparison of state-of-the-art classification models against their TTA and BayTTA counterparts, in terms of accuracy (%) and STD, after their application on the skin cancer, breast cancer, and chest X-ray datasets considered in our study. The highest overall accuracy per dataset is highlighted in bold.
@@
ARL ([61]) 86.21 $\pm$ 0.42 84.87 $\pm$ 0.15 64.58 $\pm$ 0.15
COVID-19 ([43]) 85.65 $\pm$ 0.61 94.92 $\pm$ 0.88 84.07 $\pm$ 1.41
LoTeNet ([45]) 74.89 $\pm$ 0.72 72.63 $\pm$ 1.24 79.48 $\pm$ 0.80
IRv2+SA ([14]) 90.92 $\pm$ 0.32 95.84 $\pm$ 0.54 87.76 $\pm$ 1.20
PCXRNet ([18]) 79.70 $\pm$ 0.21 93.05 $\pm$ 0.48 79.15 $\pm$ 1.46
@@
Table 6: Comparison of state-of-the-art classification models against their TTA and BayTTA counterparts, in terms of precision (PR (%)), recall (RE (%)), and F1-score (FS (%)), after their application on the skin cancer, breast cancer, and chest X-ray datasets considered in our study. The highest overall accuracy per dataset is highlighted in bold.
@@
ARL ([61]) 86.66 85.80 86.09 89.60 83.15 84.33 81.46 51.46 41.57
COVID-19 ([43]) 84.78 83.44 84.32 94.29 94.42 94.29 80.24 98.97 88.61
LoTeNet ([45]) 75.38 86.50 66.80 75.22 62.25 68.13 77.53 96.41 85.94
IRv2+SA ([14]) 89.60 90.55 90.04 97.68 95.16 95.52 84.24 98.97 91.03
PCXRNet ([18]) 79.98 99.51 88.69 93.03 73.90 82.37 75.22 99.37 85.63
@@
4.3.1 Results for medical image datasets
@@
First, we evaluate its performance in the case when it was used in combination with five pre-trained CNN models, including VGG-16, MobileNetV2, DenseNet201, ResNet152V2, and InceptionResNetV2, to determine the most suitable backbone model for the skin cancer (Table 4), breast cancer (Table 4), and chest X-ray (Table 4) datasets. In this study, we executed the baseline models three times, computing the mean accuracy and the standard deviation (STD) of accuracy. We applied TTA during the testing phase on the baseline models to assess the model accuracy and conducted a comparative analysis with BayTTA, in which the BMA method was used for model averaging. We considered each original image and six of its random augmentations (obtained through rotations) during the testing phase, evaluating the mean accuracy and STD obtained with these random augmentations. The results reported in Tables 4 demonstrate that the use of BayTTA allowed us to outperform the baseline and TTA-based models, achieving the highest accuracy and significantly reducing the standard deviation values for the skin cancer dataset. Regarding the breast cancer dataset, BayTTA outperforms the baseline and TTA-based models in accuracy for the three out of five pre-trained CNNs, i.e. VGG-16, MobileNetV2, and ResNet152V2. Nonetheless, the BayTTA-based networks consistently exhibited the lowest STD values, as outlined in Table 4. In the case of the imbalanced chest X-ray dataset (see Table 4), the combination of TTA and BMA (i.e. BayTTA) allowed us to improve the baseline and TTA results in all cases in terms of both accuracy and STD.
Table 7: Comparison of the baseline CNN model accuracy (%) $\pm$ STD performance against their TTA and BayTTA versions on the CRISPOR dataset. The highest accuracy per column is in bold. The asterisk (*) denotes the highest overall accuracy obtained by the models.
@@
Table 8: Comparison of the baseline CNN model accuracy (%) $\pm$ STD performance against their TTA and BayTTA versions on the GUIDE-seq dataset. The highest accuracy per column is in bold. The asterisk (*) denotes the highest overall accuracy obtained by the models.
@@
Table 9: Comparison of state-of-the-art classification models against their TTA and BayTTA counterparts, in terms of accuracy (%) and STD on the CRISPOR and GUIDE-seq gene editing datasets. The highest accuracy per column is highlighted in bold.
@@
ARL ([61]) 94.20 $\pm$ 1.19 91.04 $\pm$ 0.31
COVID-19 ([43]) 99.50 $\pm$ 0.20 93.87 $\pm$ 0.96
LoTeNet ([45]) 98.85 $\pm$ 0.43 93.38 $\pm$ 0.99
IRv2+SA ([14]) 99.38 $\pm$ 0.24 91.22 $\pm$ 1.03
PCXRNet ([18]) 95.69 $\pm$ 2.51 91.42 $\pm$ 0.97
@@
Table 10: Comparison of state-of-the-art classification models against their TTA and BayTTA counterparts, in terms of precision (PR (%)), recall (RE (%)), and F1-score (FS (%)) on the CRISPOR and GUIDE-seq gene editing datasets. The highest accuracy per column is highlighted in bold.
@@
ARL ([61]) 88.33 93.66 90.72 33.91 56.81 42.38
COVID-19 ([43]) 98.03 98.07 98.04 20.73 33.30 37.93
LoTeNet ([45]) 89.74 92.10 94.59 33.30 12.12 17.74
IRv2+SA ([14]) 99.73 95.20 96.71 37.86 41.91 37.54
PCXRNet ([18]) 74.13 98.01 84.43 32.90 69.67 43.47
@@
Second, we conducted a comparative study of the TTA-based, BayTTA-based, and baseline state-of-the-art classification models which have been recently used in the literature to process medical image data (see Tables 6 and 6). To ensure a thorough performance assessment, we reported the experimental results for the four following key metrics: accuracy, precision, recall, and F1-score. Our analysis revealed that among state-of-the-art models compared, COVID-19 and IRv2+SA provided the highest overall accuracy, surpassing the results of ARL, LoTeNet, and PCXRNet models. Thus, we used the TTA and BayTTA frameworks in combination with these two top-performing models. In this experiment, we considered original images as well as ten random augmentation samples generated for each of them using rotations during the test phase. Our results, reported in Table 6, indicate that the use of TTA leads to a decrease in the accuracy of IRv2+SA, but the performance of TTA is notably worse for the imbalanced chest X-ray dataset. Nonetheless, BayTTA improves the models accuracy even in cases of suboptimal performance of TTA, demonstrating its high effectiveness. The model performances in terms of precision, recall, and F1-score (see Table 6) also demonstrate a much higher robustness of the proposed BayTTA computational framework, compared to the TTA-based and baseline models.
4.3.2 Results for gene editing datasets
Similarly to medical image data, we first evaluated the performance of the proposed BayTTA method on gene editing datasets using it in combination with four pre-trained CNN models VGG-16, MobileNetV2, DenseNet201, ResNet152V2 as well as a custom CNN model with 5 layers. A comparison against the TTA-based and baseline models was carried out. The results obtained are reported in Tables 10 and 10.

### v2

@@ IV Experimental evaluations and discussion
IV-A Data used in evaluation
We conducted our experiments on three publicly available medical image datasets: skin cancer, breast cancer, and chest X-ray data, which are freely available on the Kaggle and Mendeley data repositories (see also Fig. 1). Subsequently, we evaluated the efficacy of the proposed BayTTA method using two well-known gene editing datasets, CRISPOR and GUIDE-seq, both generated through the CRISPR-Cas-9 technology [39]. We adopted an effective one-hot encoding strategy, originally proposed by Charlier et al. [54], to represent sgRNA and DNA sequence pairs as binary images from which off-target effects can be predicted by advanced DL models (see Fig. 4).
IV-A1 Medical image datasets
@@
Chest X-ray dataset is a large collection of X-ray images extracted from a publicly available medical image database [^3] https://easy.dans.knaw.nl/ui/datasets/id/easy-dataset:102451 [55]. The dataset contains 5216 training images, among which 1341 are categorized as Normal and 3875 as Pneumonia samples. Additionally, the test dataset includes 624 images, among which 234 are classified as Normal and 390 as Pneumonia samples.
IV-A2 Gene editing datasets
CRISPOR database, organized and maintained by Maximilian Haeussler [14] aggregates widely-used public datasets aimed at quantifying on-target guide efficiency and detecting off-target cleavage sites [^4] http://crispor.tefor.net. The dataset we selected from this database (see also [54]) comprises $18,211$ black and white training images (each with $8\times 23$ pixel dimension), among which 18,112 are categorized as on-targets and 99 as off-targets. Additionally, the dataset includes 7806 testing images of the same dimension, 7763 of which are classified as on-targets and 43 as off-targets.
GUIDE-seq is one of the pioneering off-target data repositories, derived from the outcomes of the GUIDE-seq technique developed by Tsai et al. [56]. It serves as an accurate framework for genome-wide identification of off-target effects. The sgRNAs used in GUIDE-seq target the following sites: VEGFA site 1, VEGFA site 2, VEGFA site 3, FANCF, HEK293 site 2, HEK293 site 3, and HEK293 site 4, wherein 28 off-targets with a minimum modification frequency of $0.1$ were identified (among 403 potential off-targets). This dataset consists of black and white images with $8\times 23$ pixel dimension. It comprises 309 training images, including 291 on-target and 18 off-target samples. Additionally, it comprises 133 testing images, including 121 on-target and 12 off-target samples.
Figure 4: Examples of visualizing CRISPR-Cas9 sgRNA-DNA sequence pairs encoded onto 8×23 matrices, then transformed into black and white images from the (a) CRISPOR and (b) GUIDE-seq gene editing datasets, respectively [54]. These images can be processed by neural networks to predict off-targets generated by CRISPR-Cas9 technology.
TABLE I: Hyperparameter configuration of the pre-trained deep learning models (VGG-16, MobileNetV2, DenseNet201, ResNet152V2, and InceptionResNetV2) for the Skin Cancer, Breast Cancer, and Chest X-ray medical image datasets considered in our study.
@@ IV-A2 Gene editing datasets
IV-B Implementation details and model settings
@@
Pre-trained CNN models. During the training phase of our first experiment, we employed five well-established convolutional networks (CNNs), namely VGG-16, MobileNetV2, DenseNet201, ResNet152V2, and InceptionResNetV2, as our feature extraction backbone, initializing weights through pre-training on the ImageNet dataset. The baseline versions of these models were initially used (without applying TTA or BayTTA), followed by their application in combination with TTA and BayTTA. All medical images were resized to the $224\times 224$ pixel size to ensure uniformity in model inputs. The hyperparameters, including the optimizer (Opt), the learning rate (LR), the batch size (BS), and the number of epochs for each of the five pre-trained models and each medical image dataset are summarized in Table I. Additionally, we employed the early stopping technique to handle overfitting [57]. During our network training, the data augmentation techniques were implemented using the following parameters: rotation range = 20, width shift range = 0.2, height shift range = 0.2, shear range = 0.2, zoom range = 0.2, horizontal flip, and vertical flip.
@@
Four pre-trained CNN models, VGG-16, MobileNetV2, DenseNet201, and ResNet152V2, as well as a custom CNN with five layers, comprising two convolutional and three fully connected dense layers, were used to predict off-targets in highly imbalanced gene editing datasets. The models were trained using the RMSprop optimizer, with a learning rate of $0.001$ and a batch size of $64$. To address potential overfitting issues a set of callback functions was used during training, specifically the ReduceLROnPlateau and EarlyStopping functions.
@@
State-of-the-art DL models intended for medical image analysis. To evaluate the performance of BayTTA, we also compared its performance against TTA and baseline, considering some state-of-the-art DL models recently used in medical image analysis, namely Attention Residual Learning (ARL) [58], COVID-19 [59], LoTeNet [60], IRv2+SA [61], and PCXRNet [62].
@@
IRv2+SA [61] explores the efficacy of soft attention in deep neural architectures. The core objective of this technique is to emphasize the significance of essential features, while mitigating the influence of noise-inducing ones. The ARL model, proposed by Zhang et al. [58], is intended for classifying skin lesions in dermoscopy images. The model architecture includes several ARL blocks, a global average pooling layer, and a classification layer. Each ARL block employs a combination of residual learning and an innovative attention-learning mechanism to boost the capacity of discriminative representations. Ozturk et al. [59] proposed a novel model based on the DarkNet method designed for automated detection of COVID-19 cases using chest X-ray images to deliver precise diagnostic results in the framework of binary and multi-class classifications. PCXRNet [62] is an attention-driven convolutional neural network designed for pneumonia diagnosis based on chest X-ray image analysis. PCXRNet incorporates a novel Condensed Attention Module (CDSE) to harness the information within the feature map channels. The LoTeNet (Locally order-less Tensor Network) model [60] is based on the use of tensor networks, a crucial tool for physicists to analyze complex quantum many-body systems.
IV-C Experimental results
@@
TABLE II: Comparison of the baseline CNN model accuracy (%) $\pm$ STD performance against the TTA and BayTTA versions on the skin cancer dataset. The highest accuracy per column is in bold. The asterisk (*) denotes the highest overall accuracy obtained by the models.
@@
TABLE III: Comparison of the baseline CNN model accuracy (%) $\pm$ STD performance against the TTA and BayTTA versions on the breast cancer dataset. The highest accuracy per column is in bold. The asterisk (*) denotes the highest overall accuracy obtained by the models.
@@
TABLE IV: Comparison of the baseline CNN model accuracy (%) $\pm$ STD performance against the TTA and BayTTA versions on the chest X-ray dataset. The highest accuracy per column is in bold. The asterisk (*) denotes the highest overall accuracy obtained by the models.
@@
TABLE V: Comparison of state-of-the-art classification models against their TTA and BayTTA counterparts, in terms of accuracy (%) and STD, after their application on the skin cancer, breast cancer, and chest X-ray datasets considered in our study. The highest overall accuracy per dataset is highlighted in bold.
@@
ARL [58] 86.21 $\pm$ 0.42 84.87 $\pm$ 0.15 64.58 $\pm$ 0.15
COVID-19 [59] 85.65 $\pm$ 0.61 94.92 $\pm$ 0.88 84.07 $\pm$ 1.41
LoTeNet [60] 74.89 $\pm$ 0.72 72.63 $\pm$ 1.24 79.48 $\pm$ 0.80
IRv2+SA [61] 90.92 $\pm$ 0.32 95.84 $\pm$ 0.54 87.76 $\pm$ 1.20
PCXRNet [62] 79.70 $\pm$ 0.21 93.05 $\pm$ 0.48 79.15 $\pm$ 1.46
@@
TABLE VI: Comparison of state-of-the-art classification models against their TTA and BayTTA counterparts, in terms of precision (PR (%)), recall (RE (%)), and F1-score (FS (%)), after their application on the skin cancer, breast cancer, and chest X-ray datasets considered in our study. The highest overall accuracy per dataset is highlighted in bold.
@@
ARL [58] 86.66 85.80 86.09 89.60 83.15 84.33 81.46 51.46 41.57
COVID-19 [59] 84.78 83.44 84.32 94.29 94.42 94.29 80.24 98.97 88.61
LoTeNet [60] 75.38 86.50 66.80 75.22 62.25 68.13 77.53 96.41 85.94
IRv2+SA [61] 89.60 90.55 90.04 97.68 95.16 95.52 84.24 98.97 91.03
PCXRNet [62] 79.98 99.51 88.69 93.03 73.90 82.37 75.22 99.37 85.63
@@
IV-C1 Results for medical image datasets
@@
First, we evaluate its performance in the case when it was used in combination with five pre-trained CNN models, including VGG-16, MobileNetV2, DenseNet201, ResNet152V2, and InceptionResNetV2, to determine the most suitable backbone model for the skin cancer (Table IV), breast cancer (Table IV), and chest X-ray (Table IV) datasets. In this study, we executed the baseline models three times, computing the mean accuracy and the standard deviation (STD) of accuracy. We applied TTA during the testing phase on the baseline models to assess the model accuracy and conducted a comparative analysis with BayTTA, in which the BMA method was used for model averaging. We considered each original image and six of its random augmentations (obtained through rotations) during the testing phase, evaluating the mean accuracy and STD obtained with these random augmentations. The results reported in Tables IV demonstrate that the use of BayTTA allowed us to outperform the baseline and TTA-based models, achieving the highest accuracy and significantly reducing the standard deviation values for the skin cancer dataset. Regarding the breast cancer dataset, BayTTA outperforms the baseline and TTA-based models in accuracy for the three out of five pre-trained CNNs, i.e. VGG-16, MobileNetV2, and ResNet152V2. Nonetheless, the BayTTA-based networks consistently exhibited the lowest STD values, as outlined in Table IV. In the case of the imbalanced chest X-ray dataset (see Table IV), the combination of TTA and BMA (i.e. BayTTA) allowed us to improve the baseline and TTA results in all cases in terms of both accuracy and STD.
TABLE VII: Comparison of the baseline CNN model accuracy (%) $\pm$ STD performance against their TTA and BayTTA versions on the CRISPOR dataset. The highest accuracy per column is in bold. The asterisk (*) denotes the highest overall accuracy obtained by the models.
@@
TABLE VIII: Comparison of the baseline CNN model accuracy (%) $\pm$ STD performance against their TTA and BayTTA versions on the GUIDE-seq dataset. The highest accuracy per column is in bold. The asterisk (*) denotes the highest overall accuracy obtained by the models.
@@
TABLE IX: Comparison of state-of-the-art classification models against their TTA and BayTTA counterparts, in terms of accuracy (%) and STD on the CRISPOR and GUIDE-seq gene editing datasets. The highest accuracy per column is highlighted in bold.
@@
ARL [58] 94.20 $\pm$ 1.19 91.04 $\pm$ 0.31
COVID-19 [59] 99.50 $\pm$ 0.20 93.87 $\pm$ 0.96
LoTeNet [60] 98.85 $\pm$ 0.43 93.38 $\pm$ 0.99
IRv2+SA [61] 99.38 $\pm$ 0.24 91.22 $\pm$ 1.03
PCXRNet [62] 95.69 $\pm$ 2.51 91.42 $\pm$ 0.97
@@
TABLE X: Comparison of state-of-the-art classification models against their TTA and BayTTA counterparts, in terms of precision (PR (%)), recall (RE (%)), and F1-score (FS (%)) on the CRISPOR and GUIDE-seq gene editing datasets. The highest accuracy per column is highlighted in bold.
@@
ARL [58] 88.33 93.66 90.72 33.91 56.81 42.38
COVID-19 [59] 98.03 98.07 98.04 20.73 33.30 37.93
LoTeNet [60] 89.74 92.10 94.59 33.30 12.12 17.74
IRv2+SA [61] 99.73 95.20 96.71 37.86 41.91 37.54
PCXRNet [62] 74.13 98.01 84.43 32.90 69.67 43.47
@@
Second, we conducted a comparative study of the TTA-based, BayTTA-based, and baseline state-of-the-art classification models which have been recently used in the literature to process medical image data (see Tables VI and VI). To ensure a thorough performance assessment, we reported the experimental results for the four following key metrics: accuracy, precision, recall, and F1-score. Our analysis revealed that among state-of-the-art models compared, COVID-19 and IRv2+SA provided the highest overall accuracy, surpassing the results of ARL, LoTeNet, and PCXRNet models. Thus, we used the TTA and BayTTA frameworks in combination with these two top-performing models. In this experiment, we considered original images as well as ten random augmentation samples generated for each of them using rotations during the test phase. Our results, reported in Table VI, indicate that the use of TTA leads to a decrease in the accuracy of IRv2+SA, but the performance of TTA is notably worse for the imbalanced chest X-ray dataset. Nonetheless, BayTTA improves the models accuracy even in cases of suboptimal performance of TTA, demonstrating its high effectiveness. The model performances in terms of precision, recall, and F1-score (see Table VI) also demonstrate a much higher robustness of the proposed BayTTA computational framework, compared to the TTA-based and baseline models.
IV-C2 Results for gene editing datasets
Similarly to medical image data, we first evaluated the performance of the proposed BayTTA method on gene editing datasets using it in combination with four pre-trained CNN models VGG-16, MobileNetV2, DenseNet201, ResNet152V2 as well as a custom CNN model with 5 layers. A comparison against the TTA-based and baseline models was carried out. The results obtained are reported in Tables X and X.